# Christian Röver, Tim Friede

### v1

@@ 1 Introduction
When evidence from a single experiment, such as a clinical trial, is not sufficiently conclusive on its own, it may sometimes help to view the data in the context of related instances (similar experiments) in order to yield more confident conclusions. This idea is explicitly implemented in shrinkage estimation, where a hierarchical model is set up accounting for estimation uncertainly at the study level as well as for variability (and similarity) between studies 1, 2; models of this kind are commonly also used in the context of meta-analysis 3, 4. The borrowing-of-information taking place between the study of primary interest and the external data may be viewed in terms of the overarching joint model as a meta-analytic-combined (MAC) approach, or, equivalently, by formulating the meta-analytic predictive (MAP) prior that explicates the information contributed by the external data to the shrinkage estimate 5. A particular special case is given when a single (“target”) study is supported by a single (“source”) study; such situations are not uncommon, and shrinkage estimation here has proven useful 6, 7, 8. Application of a hierarchical model makes it behave dynamically in the sense that more or less information is borrowed, depending on the apparent similarity of target and source data 6. When considering this case in terms of the implied MAP prior, the “meta-analysis” involved here is based on a single study, which may appear somewhat counterintuitive at first. However, such a smallest possible meta-analysis does not pose a conceptual problem; it only implies that, due to the particular sparsity of data, prior specification within the model receives special importance, a problem which, however, is common in meta-analysis of few studies in general 9], and which analogously applies for alternative borrowing methods, such as power priors or bias allowance models [8, 10. Since others appear to have struggled with or shied away from the idea of a single-study meta-analysis 11, 12, it seems worth while investigating this special case a bit closer.
@@ 2.1 The normal-normal hierarchical model (NNHM)
Often the overall mean $\mu$ is the aim of the analysis, while sometimes the study-specific parameters $\theta_{i}$ are also of interest 13, 6. The heterogeneity, while important, usually remains a nuisance parameter. In the context of “shrinkage estimation” of the $\theta_{i}$, an interesting aspect is that the problem may be motivated in two ways; classically, one may think of shrinkage estimation as a joint analysis of all ($k$) estimates, which also returns estimates of any $\theta_{i}$ parameter along the way; this is also denoted as the meta-analytic-combined (MAC) approach. The problem, however, may also be factored into the evidence stemming from the $i$th study alone, as well as the information provided by the remaining ($k-1$) estimates. Shrinkage estimation then may be interpreted as the analysis of the $i$th study, based on a prior distribution that results as the predictive distribution derived from a meta-analysis of the other ($k-1$) studies; this prior is denoted as the meta-analytic-predictive (MAP) prior. Both MAC and MAP approaches are equivalend and yield identical shrinkage estimates 5.
In the following, we will focus on the special case of only two studies ($k=2$). For the shrinkage estimate ($\theta_{2}$), this implies a MAP prior that is based on a single study (i.e., the data provided through $y_{1}$ and $s_{1}$). While this may appear odd at first, the idea readily applies also in this special case, as will be demonstrated in the following. Analysis may generally be performed based on informative or uninformative priors for $\mu$, while a proper, informative prior is required for $\tau$ 4, 9. The case of only $k=2$ studies is also closely connected to the related concepts of power prior 6 or bias allowance models 14, 10.
2.2 Uniform effect prior
Priors for the overall mean parameter ($\mu$) in the NNHM may be specified as informative or as uninformative. For (more or less informative) priors, normal distributions are an obvious choice, also since these lead to analytically simple inference. Quite commonly, effect priors however are chosen as uninformative and (improper) uniform, not least due to certain analogies to frequentist meta-analysis procedures 4. In case of an (improper, non-informative) uniform prior for $\mu$, certain expressions turn out particularly simple, which is why we will focus on this particular, yet insightful, common and practically relevant case in the following. As the (improper) uniform prior constitutes the limiting case of an increasingly uninformative effect prior, the following considerations may also be viewed as relating to the limiting behaviour for increasingly uninformative priors (e.g., for normal priors when their variance approaches infinity). The uniform effect prior leads to a normal conditional posterior for the overall mean $\mu$, with moments given by
@@ 2.2 Uniform effect prior
and a marginal heterogeneity likelihood that is constant (independent of $\tau$), so that the heterogeneity’s posterior equals its prior 4. This seems reasonable, since (as long as the effect prior is uniform) a single observation $y_{1}$ does not provide information on the heterogeneity $\tau$.
@@ 2.3 The MAP prior for the effect in a new study
The MAP prior results as the posterior predictive distribution for a “new” (second) study’s effect (here: $\theta_{2}$) given the data from the first study ($y_{1}$, $s_{1}$). In the NNHM framework, the conditional predictive distribution again is normal with mean
@@
(see also (3) and (2)). These expressions make sense in the present context: We know $\theta_{1}$ with accuracy given by the standard error $s_{1}$, and we know that the difference between $\theta_{1}$ and $\theta_{2}$ is normally distributed with variance $2\tau^{2}$, so that the (conditional) variance expression results as a corresponding sum 6. As pointed out above, information on heterogeneity ($\tau$) is based on the prior only; the heterogeneity in this context generally requires a proper, informative prior (since $k<3$) 4, 9.
The eventual (marginal) predictive distribution (marginalized over the distribution of $\tau$) hence results as a normal scale mixture 15, 16 with fixed mean (4) and with variance as given in (5), where $\tau$ is distributed according to the specified prior. In particular, one may think of the MAP prior as the first study’s point estimate $p(\theta_{1}|y_{1},s_{1})$ convolved with the prior predictive distribution $p(\theta_{2}|\theta_{1},\tau)$ and then marginalized over $\tau$. The MAP prior is symmetric around $y_{1}$, and its (marginal) variance results from (5) as
@@
where the expectation $\mathrm{E}[\tau^{2}]$ depends on the assigned heterogeneity prior. For a range of common prior specifications, this expectation may be derived analytically; Table 3 in Appendix A.2 lists some popular cases. From this expression, one can see that the relative magnitudes of $s_{1}^{2}$ and $\mathrm{E}[\tau^{2}]$ determine whether the resulting MAP prior’s variance is dominated by estimation uncertainty (regarding $\theta_{1}$) or anticipated heterogeneity ($\tau$). These two variance components may in fact also be considered to reflect so-called “type A” and “type B” uncertainties relating to measurement uncertainty and background knowledge, respectively, which together sum up to form the combined standard uncertainty $u_{c}$ (the square root of (6)) 17, 18.
Since the MAP prior results as a normal scale mixture, it is generally heavier-tailed than a normal distribution, which has implications for the resulting operating characteristics. A heavy-tailed (MAP-) prior means that in combination with a (“shorter-tailed”) normal likelihood, the likelihood will dominate in case of a prior-data conflict 19. Such robustness properties have in fact been noted and demonstrated in the meta-analysis context, as these lead to a dynamic borrowing behaviour 6. The MAP priors’ tail behaviour will also be illustrated for some examples below (Figures 4 and 5).
Another way to quantify the precision of a prior distribution is by translating it into a number of observations (effective sample size, ESS) that would in a certain sense convey an equivalent amount of information; in the following, we will quote ESSs based on the expected local-information ratio ($\mathrm{ESS}_{\mathrm{ELIR}}$) 20.
@@ 2.4 The bias allowance model connection
In the 2-study case, there is a one-to-one correspondence between the NNHM and a simple bias allowance model 10; instead of the NNHM assumption (2) in combination with a uniform effect prior and heterogeneity prior $p_{\star}(\tau)$ as in Section 2.1, one may specify
@@
with prior $p(\beta)=\frac{1}{\sqrt{2}}\,p_{\star}\bigl{(}\frac{\beta}{\sqrt{2}}\bigr{)}$ for the standard deviation $\beta$ 6. This “reference model” is different in that one estimate (the reference, or “target” $y_{2}$) directly relates to $\alpha$, while the other one (the “source” $y_{1}$) is associated with an additional offset to account for potential bias. The shrinkage estimates of $\theta_{i}$, however, can be shown to be identical in both models as long as a uniform prior for the effect ($\mu$) is used 6. The reference model may be considered a variation of Pocock’s bias model or, more generally, a bias allowance model 21, 14, 10. The $\beta$ parameter, which only differs from $\tau$ by a scaling factor of $\sqrt{2}$, may also help motivating a heterogeneity prior, as it directly relates to the expected difference between $\theta_{2}$ and $\theta_{1}$, without reference to a common overall mean $\mu$.
@@ 2.5 The power prior connection
There is also is a connection to a so-called power prior, which has been proposed as an approach for deliberate down-weighting of prior information. It is intended for a prior distribution that itself results as a posterior, and the power prior results from applying an exponent $a_{0}$ (with $0\leq a_{0}\leq 1$) to its likelihood contribution 22, 21. When conditioning on a fixed $\tau$ value, the (conditional) MAP prior is normal with moments given in (4) and (5); in particular, note that $\tau^{2}$ acts additively on the “plain” variance ($s_{1}^{2}$). In this context, a power prior with fixed exponent $a_{0}$ on the other hand would correspond to a $\mathrm{Normal}\Bigl{(}y_{1},\,\frac{s_{1}^{2}}{a_{0}}\Bigr{)}$ distribution, where the (inverse) exponent acts multiplicatively on the variance. Both MAP and power prior then are identical if $a_{0}=\bigl{(}2\frac{\tau^{2}}{s_{1}^{2}}+1\bigr{)}^{-1}$ 6, 23. While a prior probability distribution for $\tau$ is readily motivated with reference to the effect scale of $y_{i}$ and $\theta_{i}$ 9, specification of a fixed $\alpha_{0}$ value remains tricky, and a prior distribution may be even harder to motivate. On the other hand, through the above functional correspondence, any prior for the heterogeneity $\tau$ immediately implies a corresponding distribution for the exponent $a_{0}$; an example is illustrated in Appendix A.3.
@@ 3.1.1 Application
Gross et al. (2020) 24 performed a randomized controlled trial (RCT) in Alport syndrome to investigate the effects of ramipril, an angiotensin-converting enzyme inhibitor (ACEi). Recruitment of participants to the RCT was hampered by the rare and pediatric nature of the disease, and so the analysis of the RCT had been planned with the inclusion of observational data from an open-label arm and a natural disease cohort 25. Time to disease progression was a co-primary endpoint, and from the observational data, a hazard ratio (HR) of 0.53 [0.22, 1.29] was estimated based on 70 patients. Only 20 patients entered into the RCT, and a HR of 0.51 [0.12, 2.20] was estimated. The data are also summarized in Table 1.
Table 1: Alport example data from Gross et al. (2020) 24
@@
1 observational 70 0.53 [0.22, 1.29] $-0.635$ $0.451$
2 RCT 20 0.51 [0.12, 2.20] $-0.673$ $0.742$
The analysis was then performed by jointly considering both log-HR estimates in a NNHM, anticipating a reasonable amount of heterogeneity between them (expressed though a $\mathrm{half\mbox{-}Normal}(0.5)$ prior for $\tau$), and deriving a shrinkage estimate for the RCT effect ($\theta_{2}$). The resulting estimate then was substantially more precise than if the RCT data were considered in isolation; the HR was estimated at 0.52 [0.19, 1.39] 24.
@@
Figure 1: Illustration of MAP-prior, likelihood and (shrinkage-) posterior for the Alport example discussed in Section 3.1 24. The horizontal lines at the bottom indicate point estimates and corresponding 95% intervals
Figure 1 illustrates MAP-prior, likelihood and posterior (shrinkage estimate) for the Alport example. The MAP prior here has a mean of $y_{i}=-0.63$ and a variance of $s_{1}+2\,\mathrm{E}[\tau^{2}]=0.45^{2}+2\times 0.5^{2}=0.84^{2}$. While the observational sample size was 70 patients, the MAP prior’s effective sample size ($\mathrm{ESS}_{\mathrm{ELIR}}$) is at only 26 patients (i.e., 37% of originally 70 actual patients) 20. The eventual shrinkage interval is only 67% as wide as the original, implying a substantial “effective gain in sample size” 6; such a precision increase would otherwise have required more than doubling the sample size (by the addition of 24 extra patients).

### v2

@@ Abstract
  • MAP priors may serve as an additional motivation for related approaches (bias allowance models, power priors)
@@ 1 Introduction
The potential of clinical research is commonly limited by data sparsity issues; such problems particularly arise in the context of rare diseases, where the number of potential study subjects is small, or in pediatric indications, where ethical considerations may limit the recruitment of patients. The large variety of rare diseases still means that a sizeable proportion of the population is affected by rare diseases, posing a considerable economic burden. Even in more common indications, data sparsity problems may arise, for example, when the focus is on smaller sub-populations, or when novel treatments or standards of care emerge. In any of these cases, the careful consideration of all potentially relevant evidence available is essential 1, 2, 3. When evidence from a single experiment, such as a clinical trial, is not sufficiently conclusive on its own, it may sometimes help to view the data in the context of related instances (similar experiments) in order to yield more confident conclusions. This idea is explicitly implemented in shrinkage estimation, where a hierarchical model is set up accounting for estimation uncertainly at the study level as well as for variability (and similarity) between studies 4, 5; models of this kind are commonly also used in the context of meta-analysis 6, 7. The borrowing-of-information taking place between the study of primary interest and the external data may be viewed in terms of the overarching joint model as a meta-analytic-combined (MAC) approach, or, equivalently, by formulating the meta-analytic predictive (MAP) prior that explicates the information contributed by the external data to the shrinkage estimate 8. A particular special case is given when a single (“target”) study is supported by a single (“source”) study; such situations are not uncommon, and shrinkage estimation here has proven useful 9, 10, 11. Application of a hierarchical model makes it behave dynamically in the sense that more or less information is borrowed, depending on the apparent similarity of target and source data 9. When considering this case in terms of the implied MAP prior, the “meta-analysis” involved here is based on a single study, which may appear somewhat counterintuitive at first. It is this perceived contradiction that we aim to address here; while a meta-analysis is commonly thought of as involving larger amounts of data, we will see that a hierarchical model may essentially be fit also to a single data point, and sensible predictions may be derived. Such a smallest-possible meta-analysis does not pose a conceptual problem, and there is no reason to abandon the general concept even if the amount of historical data drops below a couple of studies. It only implies that, due to the particular sparsity of data, prior specification within the model receives special importance, a problem which, however, is common in meta-analysis of few studies in general 12, and which analogously applies for alternative (and closely related) borrowing methods, such as power priors or bias allowance models 11, 13. Since others appear to have struggled with or shied away from the idea of a single-study meta-analysis where in fact it may have been a viable option 14, 15, it seems worthwhile to investigate this special case a bit closer. Closer inspection of this particular case then also highlights how the properties of this MAP prior materialize, as well as its close connection to bias allowance and power prior approaches.
@@ 2.1 The normal-normal hierarchical model (NNHM)
Often the overall mean $\mu$ is the aim of the analysis, while sometimes the study-specific parameters $\theta_{i}$ are also of interest 16, 9. The heterogeneity, while important, usually remains a nuisance parameter. In the context of “shrinkage estimation” of the $\theta_{i}$, an interesting aspect is that the problem may be motivated in two ways; classically, one may think of shrinkage estimation as a joint analysis of all ($k$) estimates, which also returns estimates of any $\theta_{i}$ parameter along the way; this is also denoted as the meta-analytic-combined (MAC) approach. The problem, however, may also be factored into the evidence stemming from the $i$th study alone, as well as the information provided by the remaining (k1k-1) estimates. Shrinkage estimation then may be interpreted as the analysis of the $i$th study, based on a prior distribution that results as the predictive distribution derived from a meta-analysis of the other (k1k-1) studies; this prior is denoted as the meta-analytic-predictive (MAP) prior. Both MAC and MAP approaches are equivalent and yield identical shrinkage estimates 8.
In the following, we will focus on the special case of only two studies ($k=2$). For the shrinkage estimate ($\theta_{2}$), this implies a MAP prior that is based on a single study (i.e., the data provided through $y_{1}$ and $s_{1}$). While this may appear odd at first, the idea readily applies also in this special case, as will be demonstrated in the following. Analysis may generally be performed based on informative or uninformative priors for $\mu$, while a proper, informative prior is required for $\tau$ 7, 12. The case of only $k=2$ studies is also closely connected to the related concepts of power prior 9 or bias allowance models 17, 13.
2.2 Uniform prior for the overall mean effect ($\mu$)
Priors for the overall mean parameter ($\mu$) in the NNHM may be specified as informative or as uninformative. For (more or less informative) priors, normal distributions are an obvious choice, also since these lead to analytically simple inference. Quite commonly, effect priors however are chosen as uninformative and (improper) uniform, not least due to certain analogies to frequentist meta-analysis procedures 7. In case of an (improper, non-informative) uniform prior for $\mu$, certain expressions turn out particularly simple, which also is why we will focus on this particular, yet insightful, common and practically relevant case in the following. As the (improper) uniform prior constitutes the limiting case of an increasingly uninformative effect prior, the following considerations may also be viewed as relating to the limiting behaviour for increasingly uninformative priors (e.g., for normal priors when their variance approaches infinity). The uniform effect prior leads to a normal conditional posterior for the overall mean effect $\mu$, with moments given by
@@ 2.2 Uniform prior for the overall mean effect ($\mu$)
and a marginal heterogeneity likelihood that is constant (independent of $\tau$), so that the heterogeneity’s posterior equals its prior 7. This seems reasonable, since (as long as the overall mean effect prior is uniform) a single observation $y_{1}$ does not provide information on the heterogeneity $\tau$.
@@ 2.3 The MAP prior for the effect in a new study
The MAP prior results as the posterior predictive distribution for a “new”, second study’s (study-specific) effect $\theta_{2}$ given the data from the first study ($y_{1}$, $s_{1}$). In the NNHM framework, the conditional predictive distribution again is normal with mean
@@
(see also (3) and (2), and the more detailed derivation in Appendix A.1). These expressions make sense in the present context: We know $\theta_{1}$ with accuracy given by the standard error $s_{1}$, and we know that the difference between $\theta_{1}$ and $\theta_{2}$ is normally distributed with variance $2\tau^{2}$, so that the (conditional) variance expression results as a corresponding sum 9. As pointed out above, information on heterogeneity ($\tau$) so far is based on the prior only; the heterogeneity in this context generally requires a proper, informative prior (since $k<3$) 7, 12.
The eventual (marginal) predictive distribution (marginalized over the distribution of $\tau$) hence results as a normal scale mixture 18, 19 with fixed mean (4) and with variance as given in (5), where $\tau$ is distributed according to the specified prior. In particular, one may think of the MAP prior as the first study’s point estimate $p(\theta_{1}|y_{1},s_{1})$ convolved with the prior predictive distribution $p(\theta_{2}|\theta_{1},\tau)$ and then marginalized over $\tau$. The MAP prior is symmetric around $y_{1}$, and its (marginal) variance results from (5) as
@@
where the expectation $\mathrm{E}[\tau^{2}]$ depends on the assigned heterogeneity prior. For a range of common prior specifications, this expectation may be derived analytically; Table A3 in Appendix A.4 lists some popular cases. From this expression, one can see that the relative magnitudes of $s_{1}^{2}$ and $\mathrm{E}[\tau^{2}]$ determine whether the resulting MAP prior’s variance is dominated by estimation uncertainty (regarding $\theta_{1}$) or anticipated heterogeneity ($\tau$). These two variance components may in fact also be considered to reflect so-called “type A” and “type B” uncertainties relating to measurement uncertainty and background knowledge, respectively, which together sum up to form the combined standard uncertainty $u_{c}$ (the square root of (6)) 20, 21.
Since the MAP prior results as a normal scale mixture, it is generally heavier-tailed than a normal distribution, which has implications for the resulting operating characteristics. A heavy-tailed (MAP-) prior means that in combination with a (“shorter-tailed”) normal likelihood, the likelihood will dominate in case of a prior-data conflict 22. Such robustness properties have in fact been noted and demonstrated in the meta-analysis context, as these lead to a dynamic borrowing behaviour 9. The MAP priors’ tail behaviour will also be illustrated for some examples below (Figures 4 and 5).
Another way to quantify the precision of a prior distribution is by relating it to a number of observations that would in a certain sense convey an equivalent amount of information. In the following, we will use two approaches to this effect. Firstly, a prior may be assessed in terms of a corresponding absolute effective sample size (ESS) (here: number of patients), this will be done by quoting ESSs based on the expected local-information-ratio ($\mathrm{ESS}_{\mathrm{ELIR}}$). This measure is based on the prior density’s curvature and it ensures predictive consistency, that is, on expectation, the posterior’s ESS will be the sum of the prior’s $\mathrm{ESS}_{\mathrm{ELIR}}$ plus the actual sample size 23. Secondly, the added information from including the prior in a specific analysis may be expressed in terms of the (relative) gain in effective sample size 9. This is based on comparing the relative width of a confidence interval with and without considering the informative prior, and then determining by what factor the sample size would have needed to be increased to yield the same precision gain (see also Appendix A.2).
@@ 2.4 The bias allowance model connection
In the 2-study case, there is a one-to-one correspondence between the NNHM and a simple bias allowance model 13; instead of the NNHM assumption (2) in combination with a uniform prior for the overall mean effect $\mu$ and heterogeneity prior $p_{\star}(\tau)$ as in Section 2.1, one may specify
@@
with prior $p(\beta)=\frac{1}{\sqrt{2}}\,p_{\star}\bigl(\frac{\beta}{\sqrt{2}}\bigr)$ for the standard deviation $\beta$ 9. This “reference model” is different in that one estimate (the reference, or “target” $y_{2}$) directly relates to $\alpha$, while the other one (the “source” $y_{1}$) is associated with an additional offset to account for potential bias. The shrinkage estimates of $\theta_{i}$, however, can be shown to be identical in both models as long as a uniform prior for the overall mean effect ($\mu$) is used 9. The reference model may be considered a variation of Pocock’s bias model or, more generally, a bias allowance model 24, 17, 13. The $\beta$ parameter, which only differs from $\tau$ by a scaling factor of $\sqrt{2}$, may also help motivating a heterogeneity prior, as it directly relates to the expected difference between $\theta_{2}$ and $\theta_{1}$, without reference to a common overall mean $\mu$.
@@ 2.5 The power prior connection
There is also is a connection to a so-called power prior, which has been proposed as an approach for deliberate down-weighting of prior information. It is intended for a prior distribution that itself results as a posterior, and the power prior results from applying an exponent $a_{0}$ (with $0\leq a_{0}\leq 1$) to its likelihood contribution 25, 24. When conditioning on a fixed $\tau$ value, the (conditional) MAP prior is normal with moments given in (4) and (5); in particular, note that $\tau^{2}$ acts additively on the “plain” variance ($s_{1}^{2}$). In this context, a power prior with fixed exponent $a_{0}$ on the other hand would correspond to a $\mathrm{Normal}\Bigl(y_{1},\,\frac{s_{1}^{2}}{a_{0}}\Bigr)$ distribution, where the (inverse) exponent acts multiplicatively on the variance. Both MAP and power prior then are identical if $a_{0}=\bigl(2\frac{\tau^{2}}{s_{1}^{2}}+1\bigr)^{-1}$ 26, 9, 27. It is interesting to note that the relationship between $a_{0}$ and $\tau$ here depends on the the ratio $\frac{\tau}{s_{1}}$; Pawel et al. (2024) 27 point out that the exponent $a_{0}$ directly relates to the “relative” heterogeneity as expressed though the popular $I^{2}$ statistic 28, which in this case simply equals $I^{2}=\frac{\tau^{2}}{\tau^{2}+s_{1}^{2}}$. The exponent may then directly be expressed as a function of the corresponding $I^{2}$ as $a_{0}=\frac{1-I^{2}}{1+I^{2}}$. While a prior probability distribution for $\tau$ is readily motivated with reference to the effect scale of $y_{i}$ and $\theta_{i}$ 12, specification of a fixed $\alpha_{0}$ value remains tricky, and a prior distribution may be even harder to motivate, as it would implicitly relate to the $I^{2}$ scale. On the other hand, through the above functional correspondence, any prior for the heterogeneity $\tau$ immediately implies a corresponding distribution for the exponent $a_{0}$; an example is illustrated in Appendix A.5.
@@ 3.1.1 Application
Gross et al. (2020) 29 performed a randomized controlled trial (RCT) in Alport syndrome to investigate the effects of ramipril, an angiotensin-converting enzyme inhibitor (ACEi). Recruitment of participants to the RCT was hampered by the rare and pediatric nature of the disease, and so the analysis of the RCT had been planned with the inclusion of observational data from an open-label arm and a natural disease cohort 30. Time to disease progression was a co-primary endpoint, and from the observational data, a hazard ratio (HR) of 0.53 [0.22, 1.29] was estimated based on 70 patients. Only 20 patients entered into the RCT, and a HR of 0.51 [0.12, 2.20] was estimated. The data are also summarized in Table 1.
Table 1: Alport example data from Gross et al. (2020) 29
@@
1 observational 70 0.53 [0.22, 1.29] 0.635-0.635 $0.451$
2 RCT 20 0.51 [0.12, 2.20] 0.673-0.673 $0.742$
The analysis was then performed by jointly considering both log-HR estimates in a NNHM, anticipating a reasonable amount of heterogeneity between them (expressed through a $\mathrm{half\mbox{-}Normal}(0.5)$ prior for $\tau$), and deriving a shrinkage estimate for the RCT effect ($\theta_{2}$). The resulting estimate then was substantially more precise than if the RCT data were considered in isolation; the HR was estimated at 0.52 [0.19, 1.39] 29.
@@
Figure 1: Illustration of MAP-prior, likelihood and (shrinkage-) posterior for the Alport example discussed in Section 3.1 29. The horizontal lines at the bottom indicate point estimates and corresponding 95% intervals
Figure 1 illustrates MAP-prior, likelihood and posterior (shrinkage estimate) for the Alport example. The MAP prior here has a mean of $y_{i}=-0.63$ and a variance of $s_{1}^{2}+2\,\mathrm{E}[\tau^{2}]=0.45^{2}+2\times 0.5^{2}=0.84^{2}$. While the observational sample size was 70 patients, the MAP prior’s effective sample size ($\mathrm{ESS}_{\mathrm{ELIR}}$) is at only 26 patients (i.e., 37% of originally 70 actual patients) 23. The eventual shrinkage interval is only 67% as wide as the original, implying a substantial “effective gain in sample size” 9; such a precision increase would otherwise have required more than doubling the sample size (by the addition of 24 extra patients). So in this case the absolute ($\mathrm{ESS}_{\mathrm{ELIR}}$) estimate matches well the observed precision gain.